Please help me solve this problem on tangents and secants:
x^2 + y^2 = 10
y = 3x

Answer by cleomenius(959)

Since y = 3x, we can use a substitution in the first equation.
x^2 +{3x)^2 = 10
10x^2 = 10
x^2 = 1
x = 1
Substituting into the first equation, we get
1 + y^2 = 10
y^2 = 9
y = 3.
This checks with the second equation, 3 = (3)(1).

Cleomenius.